# Goodness-of-fit test for multi-layer stochastic block models

### v1

@@ 6 Numerical experiments
  • 2.
    For the critical transition regime ($K_{0}=K$), we observe that $T$ collapses near zero (bold values), with $|T|\leq 0.2570$ across all $K$. This sudden drop confirms asymptotic normality shown in Theorem 1.
  • 3.
@@
The consistency of this pattern: steep divergence for $K_{0}<K$, immediate normalization at $K_{0}=K$, and sustained stability for $K_{0}>K$, validates the efficacy of the sequential testing algorithm NAST in identifying the true community count.
@@ 7 Real data
In Figure 2, we plot the absolute value of the test statistic $|T|$ against increasing $K_{0}$ for these real networks. The result reveals that $|T|$ computed across various candidate community counts $K_{0}$ frequently exceeds standard critical values significantly—for instance, $z_{1-\alpha/2}\approx 1.96$ at $\alpha=0.05$ or $z_{1-\alpha/2}\approx 4.5648$ at $\alpha=0.000005$. This deviation from the asymptotic normality under $H_{0}$ predicted by Theorem 1 and observed in simulated settings of Table 3 suggests potential model misspecification in real data. To robustly estimate the true number of communities $K$ under these conditions, we propose an alternative strategy leveraging the relative change in $|T|$ across sequential $K_{0}$ values.
The estimation procedure is as follows:
  • 1.
    For a given multi-layer network, compute $T(K_{0})$ for $K_{0}=1,2,\ldots,K_{\text{max}}$. We set $K_{\text{max}}=\lceil\sqrt{n}\rceil$, reflecting our assumption that the true $K$ in real-world networks is typically moderate, where $\lceil x\rceil$ denotes the smallest integer that is greater than or equal to $x$.
  • 2.
    Record the sequence of absolute values $|T(1)|,|T(2)|,\ldots,|T(K_{\text{max}})|$.
  • 3.
    Define the ratio:
    $\eta_{K_{0}}=\frac{|T(K_{0}-1)|}{|T(K_{0})|}\quad\text{for}\quad K_{0}=2,3,\ldots,K_{\text{max}}.$
  • 4.
    Estimate the number of communities $\hat{K}$ as:
    $\hat{K}=\underset{K_{0}=2,3,\ldots,K_{\text{max}}}{\text{argmax}}\eta_{K_{0}},$
    where we assume $K\geq 2$ for real networks, as $K=1$ precludes meaningful community structure analysis.
This approach identifies $K_{0}$ where $|T|$ exhibits the largest relative drop compared to $K_{0}-1$, corresponding to the transition from underfitting $K_{0}<K$ to adequate fitting $K_{0}=K$ observed in Table 3. Crucially, it does not rely on the asymptotic $N(0,1)$ distribution by exploiting the characteristic phase-transition behavior of $T$:
  • 1.
    For the underfitting case when $K_{0}<K$, $|T|$ remains large due to unmodeled community structure (leading to $H_{0}$ rejection).
  • 2.
    For the critical fitting case when ($K_{0}=K$), $|T|$ collapses sharply (leading to $H_{0}$ acceptance).
  • 3.
    For the overfitting case when ($K_{0}>K$), $|T|$ stabilizes near zero for $K_{0}$ slightly larger than $K$.
Maximizing $\eta_{K_{0}}$ detects this collapse point. This aligns with the pattern observed in Table 3, where $\eta_{K}$ peaks dramatically at the true $K$ (e.g., $\eta_{2}\approx 332.5449/0.2570\approx 1294$, $\eta_{5}\approx 17.2351/0.0081\approx 2128$). This method provides a robust empirical alternative for estimating $K$ when the formal hypothesis test exhibits size distortion in real-world networks.
@@
Based on the empirical results presented in Figure 3, the proposed ratio-based estimator $\eta_{K_{0}}$ can identify the number of communities across all eight real-world networks by detecting the sharpest relative drop in the absolute test statistic $|T|$. Based on this figure, we have:
  • 1.
    Our method exactly determines the true $K$ for all four single-layer networks with known ground truth $K$. For Dolphins, Football, Polbooks, and UKfaclty, $\eta_{K_{0}}$ peaks at $K_{0}=2,11,2$, and $3$, respectively, aligning with their true number of communities.
  • 2.
    For the four real multi-layer networks with unknown true $K$, our method identifies a proper number of communities for them. In detail, our method estimates the number of communities for Lazega Law Firm, C.Elegans, CS-Aarhus, and FAO-trade as 2, 2, 5, and 3, respectively.
@@ 8 Conclusion
This paper introduces a principled framework for determining the number of communities in the multi-layer stochastic block model, addressing a critical gap in the analysis of complex multi-layer networks. Our approach centers on a novel spectral-based goodness-of-fit test leveraging a normalized aggregation of layer-wise adjacency matrices. Under mild regularity conditions, we establish the asymptotic normality of a test statistic derived from the trace of the cubed normalized matrix when the candidate community count $K_{0}$ is correct. This theoretical foundation facilitates a computationally efficient sequential testing procedure, which iteratively evaluates increasing values of $K_{0}$ until the null hypothesis $H_{0}:K=K_{0}$ is accepted. Numerical experiments on both simulated and real-world multi-layer networks demonstrate the accuracy and efficiency of our method in recovering the true number of communities. To the best of our knowledge, this is the first method for determining $K$ in multi-layer networks with rigorous theoretical guarantees.
@@
CRediT authorship contribution statement
Huan Qing: Conceptualization, Data curation, Formal analysis, Funding acquisition, Methodology, Software, Visualization, Writing – original draft, Writing - review &\& editing.
Declaration of competing interest
The authors declare no competing interests.
Data availability
Data will be made available on request.
Acknowledgements
Huan Qing was supported by the Scientific Research Foundation of Chongqing University of Technology (Grant No. 2024ZDR003) and the Science and Technology Research Program of Chongqing Municipal Education Commission (Grant No. KJQN202401168).

### v2

@@ 3 NAST 0 1 1 1 1 1 1
  • For the critical transition regime ($K_{0}=K$), we observe that $T$ collapses near zero (bold values), with $|T|\leq 0.0947$ across all $K$. This sudden drop confirms asymptotic normality shown in Theorems 1 and 2.
@@
The consistency of this pattern in the test statistic $T$: steep divergence for $K_{0}<K$, immediate normalization at $K_{0}=K$, and sustained stability for $K_{0}>K$, validates the efficacy of the sequential testing algorithm NAST in identifying the true community count.
Table 4: Average values of the ratio statistic $\eta_{K_{0}}$ over 200 Monte Carlo trials for different candidate community numbers $K_{0}$ when the true number of communities $K\in\{2,3,4,5,6\}$. The first $\eta_{K_{0}}$ value satisfying $\eta_{K_{0}}>t_{\text{ratio},n}$ with $t_{\text{ratio},n}=\log n\approx 6.9078$ when $n=1000$ for each true $K$ is highlighted in bold. Note that $\eta_{K_{0}}$ is defined from $K_{0}=2$, while $K_{0}=1$ is recorded in Table 3. The settings of network parameters $(n,L,\theta,\{B_{\ell}\}^{L}_{\ell=1})$ are the same as Experiment 3, with the sparsity parameter $\rho$ being $0.1$.
$K_{0}$
2 3 4 5 6 7 8 9 10
$K=2$ 1629.01 1.10 1.66 1.71 1.49 0.91 0.85 0.96 0.95
$K=3$ 2.62 521.08 1.09 1.03 1.03 1.05 1.46 1.57 2.40
$K=4$ 1.80 2.59 132.78 1.01 1.09 1.00 1.00 1.08 1.10
$K=5$ 1.53 1.79 2.55 63.43 1.00 1.10 1.04 1.27 1.03
$K=6$ 1.40 1.52 1.78 2.52 53.66 1.05 0.99 1.03 1.33
Table 4 provides decisive empirical validation of the theoretical phase transition of the ratio statistic $\eta_{K_{0}}$, as established in Theorem 4. The data exhibit a clear and consistent pattern across all true community counts $K$:
  • Underfitting regime ($K_{0}<K$): For every $K\in\{2,3,4,5,6\}$, the values of $\eta_{K_{0}}$ remain stochastically bounded, with all entries satisfying $\eta_{K_{0}}\leq 2.62$. This observation confirms the theoretical prediction $\eta_{K_{0}}=O_{P}(1)$ for $K_{0}<K$.
  • Critical transition ($K_{0}=K$): A distinct and dramatic peak in $\eta_{K_{0}}$ occurs precisely at $K_{0}=K$ (bold values). The statistic jumps to values orders of magnitude larger than those in the underfitting regime (e.g., $\eta_{4}=132.78$ for $K=4$, compared to $\eta_{3}=2.59$). This sharp increase empirically demonstrates the divergence in probability $\eta_{K}\xrightarrow{P}\infty$, driven by the diverging numerator |T(K1)||T(K-1)| under $H_{1}$ and the $O_{P}(1)$ denominator $|T(K)|$ under $H_{0}$.
  • Overfitting regime ($K_{0}>K$): For $K_{0}>K$, $\eta_{K_{0}}$ returns to magnitudes comparable to the underfitting regime, reflecting the absence of further structural signal beyond the true model order.
The empirical dichotomy—boundedness for all underspecified models versus a pronounced, unique peak at the true model—robustly underpins the stopping rule of the SR-NAST algorithm. By sequentially testing until $\eta_{K_{0}}$ surpasses a slowly growing threshold (e.g., $t_{\mathrm{ratio},n}=\log n$), the algorithm consistently and efficiently identifies the true $K$, leveraging this theoretically guaranteed phase transition. The consistency of this pattern across all tested values of $K$ confirms the practical reliability of the ratio-based approach for sequential model selection.
@@ 7 Real data
[…]
@@
  • SR-NAST exactly determines the true $K$ for all four single-layer networks with known ground truth $K$. For Dolphins, Football, Polbooks, and UKfaclty, $\eta_{K_{0}}$ peaks at $K_{0}=2,11,2$, and $3$, respectively, aligning with their true number of communities.
  • For the four real multi-layer networks with unknown true $K$, SR-NAST identifies a proper number of communities for them. In detail, SR-NAST estimates the number of communities for Lazega Law Firm, C.Elegans, CS-Aarhus, and FAO-trade as 2, 2, 5, and 3, respectively.
@@ 8 Conclusion
This paper introduces a principled framework for determining the number of communities in the multi-layer stochastic block model, addressing a critical gap in the analysis of complex multi-layer networks. Our approaches centers on a novel spectral-based goodness-of-fit test leveraging a normalized aggregation of layer-wise adjacency matrices. Under mild regularity conditions, we establish the asymptotic normality of a test statistic derived from the trace of the cubed normalized matrix when the candidate community count $K_{0}$ is correct; we also build the asymptotic power of this test statistic under the alternative hypothesis. The two theoretical foundations facilitate two computationally efficient sequential testing procedures, where both algorithms iteratively evaluate increasing values of $K_{0}$ until the null hypothesis $H_{0}:K=K_{0}$ is accepted. Numerical experiments on both simulated and real-world multi-layer networks demonstrate the accuracy and efficiency of our methods in recovering the true number of communities. To the best of our knowledge, this is the first work for determining $K$ in multi-layer networks with rigorous theoretical guarantees.
@@
[…]
Declarations
Conflict of interest The author declares no conflict of interest.